I need help in giving an exact and approximate solution to three decimal places. . x^2+6x+9=81
Answer by checkley75(3666)

x^2+6x+9=81
x^2+6x+9-81=0
x^2+6x-72=0
(x+12)(x-6)=0
x+12=0
x=-12 ans.
x-6=0
x=6 ans.